OpenStudy (anonymous):

solve the equation for 0

OpenStudy (mani_jha):

Now substitute the value of cos45=1/sqrt(2)

OpenStudy (anonymous):

how did you get to that?

OpenStudy (anonymous):

sorry, im still confused

OpenStudy (anonymous):

oh! so it comes to be 2cosx (root2/2)=root2

OpenStudy (mani_jha):

The value of cos45 is a standard value. You should know it. \[2cosxcos45=2cosx(1/\sqrt{2})=\sqrt{2}cosx\] Now proceed

OpenStudy (anonymous):

ohh. so it ends up being cosx=root2/root2 so cosx=1?

OpenStudy (mani_jha):

Yes. Now, do you know how to find the values for x?

OpenStudy (anonymous):

0 and 360 degrees?

OpenStudy (mani_jha):

Yes. The general solution of cosx=1 is x=2nPI

OpenStudy (anonymous):

thank you!